A car travels 13 km in 2 hours and 2 km in 1 hour







in the same direction. What is the car's average







speed in km/hr and m/s for the whole journey?

1
Expert's answer
2022-03-08T08:56:34-0500

The average speed

"v_{ave}=\frac{d_1+d_2}{t_1+t_2}=\frac{13+2}{2+1}\\
=5\:\rm km/hr=18\: m/s"